OpenStudy (anonymous):

find the limit as x -> -2 (x^2+1)/(3x^2-2x+5)

OpenStudy (anonymous):

replace \(x\) by \(-2\) and see what you get

OpenStudy (anonymous):

i got like -1/7? i just don't know if it's right :(

OpenStudy (anonymous):

\((-2)^2+1=4+1=5\) for the numerator

OpenStudy (anonymous):

\[3(-2)^2-2\times (-1)+5\] is your denominator i doubt you get \(-\frac{1}{7}\)

OpenStudy (anonymous):

5/21?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

in fact if my arithmetic is correct, the denominator is \(19\) making the limit \(\frac{5}{19}\)

OpenStudy (anonymous):

course my arithmetic could be wrong

OpenStudy (anonymous):

check it

OpenStudy (anonymous):

isnt the equation 2x -2?

OpenStudy (anonymous):

lol you are right, i am wrong

OpenStudy (anonymous):

not 2x -1?

OpenStudy (anonymous):

yeah i made a stupid mistake you are correct \(\frac{5}{21}\)

OpenStudy (anonymous):

es otay, (: do you know how to find a fourth root on a calculator? o: :(

OpenStudy (anonymous):

depends on the calculator you could find the square root twice if it is not a fancy one

OpenStudy (anonymous):

or just use this http://www.wolframalpha.com/input/?i=fourth+root%2816%29

OpenStudy (anonymous):

ti-84? o:

OpenStudy (anonymous):

i don't really use calculators when i am on line wolfram does it all, and is very very easy to use see the link i sent